A conducting loop is made in the form of two squares of sides s1 = 3cm and s2 = 6.3 cm as shown. At time t = 0, the loop enters a region of length L = 17.3 cm that contains a uniform magnetic field B = 1.2 T, directed in the positive z-direction. The loop continues through the region with constant speed v = 35 cm/s. The resistance of the loop is R = 1.7 Ω.

1)At time t = t1 = 0.029 s, what is I1, the induced current in the loop? I1 is defined to be positive if it is in the counterclockwise direction.

2)At time t = t2 = 0.64 s, what is I2, the induced current in the loop? I2 is defined to be positive if it is in the counterclockwise direction.

3)What is Fx(t2), the x-component of the force that must be applied to the loop to maintain its constant velocity v = 35 cm/s at t = t2 = 0.64 s?

4)At time t = t3 = 0.523 s, what is I3, the induced current in the loop? I3 is defined to be positive if it is in the counterclockwise direction.

5)Consider the two cases shown above. How does II, the magnitude of the induced current in Case I, compare to III, the magnitude of the induced current in Case II? Assume s2 = 3s1.

h16_constantBD

II < III

II = III

II > III
Answer #1

Given , Side of square 1 , s1 = 3 cm = 0.03 m { 1 m = 100 cm }

Side of square 2 , s2 = 6.3 cm = 0.063 m

Magnetic field , B = 1.2 T

Length of region of magnetic field , L = 17.3 cm = 0.173 m

Speed of loop = 35 cm/s = 0.35 m/s

Resistance of loop , R = 1.7 \Omega

( a ) The induced current , I1 at time , t1 = 0.029 second can be calculated as :

At t1 = 0.029 s , distance of loop in magnetic field , d = v * t1 = 0.035 * 0.029 = 0.0101 m = 1.01 cm

\varepsilon = -\frac{d\phi }{dt}

where \varepsilon = Induced EMF or voltage

\phi = flux

\varepsilon_{1} = -\frac{d\phi }{dt} = B.v.s_{1}

\varepsilon_{1} = - 1.2*0.35*0.03

\varepsilon_{1} = - 0.0126 V { negative due to the increasing \phi }

Now , \varepsilon _{1}=I_{1}R

I_{1}=\frac{\varepsilon_{1} }{R}

I_{1}=\frac{-0.0126 }{1.7}=-0.0074 A=-7.4mA

Hence , the induced current in the loop at t1 = 0.029 second is - 0.0074 A or - 7.4 mA

( b ) The induced current , I2 at time , t2 = 0.64 second can be calculated as :

At t2 = 0.64 s , distance of loop in magnetic field , d = v * t2 = 0.35 * 0.64 = 0.224 m

distance of loop outside field = 0.224 - 0.173 = 0.051 m

{ This is greater than s1, so we must be looking at an area of s2 }

\varepsilon_{2} = -\frac{d\phi }{dt} = B.v.s_{2}

\varepsilon_{2} = 1.2*0.35*0.063

\varepsilon_{2} = 0.02646V    { positive due to the decreasing \phi }

\varepsilon _{2}=I_{2}R

I_{2}=\frac{\varepsilon_{2} }{R}

I_{2}=\frac{0.02646 }{1.7}=0.0156 A=15.6mA

Hence , the induced current in the loop at t2 = 0.64 second is 0.0156 A or 15.6 mA

( c ) The x-component of force , Fx at t2 = 0.64 second can be calculated as :

Force , F is given by :

F = I.L*B

where I = current

L = length { L in this case is the length of the square in the magnetic field }

B = magnetic field

The bottom and top part of the square have forces that act in opposite directions and cancel .

The x - component of force is given by :

F_{x} = I_{2}.s_{2}*B

F_{x} =0.0156 *0.063*1.2

F_{x} =0.001179N

F_{x} =1.179*10^{-3}N

Hence , the x - component of force is 0.001179 Newton

( d ) The induced current , I3 at time , t3 = 0.523 second can be calculated as :

At t3 = 0.523 s , distance of loop in magnetic field , d = v * t3 = 0.35 * 0.523 = 0.183 m

distance of loop outside field  = 0.183 - 0.173 = 0.010 m

{ This is smaller than s1, so we must be looking at an area of s1 }

\varepsilon_{3} = -\frac{d\phi }{dt} = B.v.s_{1}

\varepsilon_{3} = 1.2*0.35*0.03

\varepsilon_{3} = 0.0126V { positive due to the decreasing \phi }

\varepsilon _{3}=I_{3}R

I_{3}=\frac{\varepsilon_{3} }{R}

I_{3}=\frac{0.0126 }{1.7}=0.0074 A=7.4mA

Hence , the induced current in the loop at t3 = 0.523 second is -0.0074 A or 7.4 mA .